OpenStudy (anonymous):

Evaluate A3, if possible. A= [7 3] [-2 0] Please help, having a hard time figuring these out

OpenStudy (jdoe0001):

A3? A * 3?

OpenStudy (jdoe0001):

\(\large A= \begin{bmatrix} 7&3\\-2&0 \end{bmatrix}\qquad A3=A\cdot 3?\)

OpenStudy (anonymous):

No its a to the 3rd power

OpenStudy (jdoe0001):

hmm all I can think of is, then \(A= \begin{bmatrix} 7&3\\-2&0 \end{bmatrix}\qquad A^3\implies \begin{bmatrix} 7&3\\-2&0 \end{bmatrix}\times \begin{bmatrix} 7&3\\-2&0 \end{bmatrix}\times \begin{bmatrix} 7&3\\-2&0 \end{bmatrix}\)

OpenStudy (anonymous):

okay thank you

OpenStudy (jdoe0001):

yw